# Hilbert’s Theorem 90, periodicity, and roots of Artin-Schreier polynomials

### v1

@@ 3 Periodicity and Hilbert’s additive Theorem 90
Surprisingly, Hilbert did not prove the additive form of Hilbert’s Theorem 90. Lang in [4, p. 215] gives a proof, which he attributes to Emil Artin. It uses the expression $R(z,y)$. By contrast Theorem 3.1(a) uses $R(y,z)$, and the form $R(y,z)=\sum_{i=0}^{n-1}x_{i}\sigma^{i}y$ with $x_{i}=\sum_{j=0}^{n-1}\sigma^{j}z$ suggests the periodicity result in parts (b,c). The next section considers closed-form expressions for the roots of Artin-Schreier polynomials.
4 Formulas for the roots of Artin-Schreier polynomials
Let $E$ be a field of prime characteristic $p$. The $p$th power map $x\mapsto x^{p}$ defines a monomorphism $E\to E$. In this section, we assume that ${\mathbb{F}}_{q}\leqslant F\leqslant E$ where $q=p^{f}$ and $E/F$ is cyclic of degree $n$ where $\textup{Gal}(E/F)=\langle\sigma\rangle$ and $\sigma(x)=x^{q}$. One such example involves finite fields $E={\mathbb{F}}_{q^{n}}$ and $F={\mathbb{F}}_{q}$, another involves function fields $E={\mathbb{F}}_{q^{n}}(T)$ and $F={\mathbb{F}}_{q}(T)$.
@@ 4 Formulas for the roots of Artin-Schreier polynomials
Suppose $E/F$ is a cyclic Galois extension of degree $n$ where $\textup{char}(F)=p$ and $\sigma(x)=x^{q}$, $q=p^{f}$, generates $\textup{Gal}(E/F)$. If $y\in E$ and $\textup{Tr}^{E}_{F}(y)=\sum_{i=0}^{n-1}y^{q^{i}}=0$, then
  1. (a)
    $x=\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}z^{q^{j}}\right)y^{q^{i}}$ is a root of $t^{q}-t-y$ where $z\in E$ satisfies $\textup{Tr}^{E}_{F}(z)=1$;
  2. (b)
  3. (c)
    the sequence $(\sum_{j=0}^{i-1}z^{q^{j}})_{i\geqslant 0}$ is periodic with period $pe$ where $e=|F(z):F|$.
In the case when $n$ is coprime to $p$, i.e. $n_{p^{\prime}}=1$, the formula above for $x$ simplifies.
@@
Consider Corollary 4.1 when $p\nmid n$, that is $n_{p}=1$. If $y\in E$ has trace $\textup{Tr}^{E}_{F}(y)=\sum_{i=0}^{n-1}y^{q^{i}}=0$, then $x=\sum_{i=0}^{n-1}\frac{i}{n}y^{q^{i}}$ satisfies $x^{q}-x-y=0$.
@@
By Corollary 4.1, we may choose $z$ to lie in $F$ and so $1=\textup{Tr}^{E}_{F}(z)=nz$. Hence $z=\frac{1}{n}$ and $x=R(y,z)=\sum_{i=0}^{n-1}\frac{i}{n}y^{q^{i}}$ satisfies $x^{q}-x-y=0$ by Theorem 4.1(a). ∎
Remark 4.3.
Corollary 4.2 shows that the roots of $t^{q}-t-y$ lie in the ${\mathbb{F}}_{q}$-subspace $\langle y,y^{q},\dots,y^{q^{n-1}}\rangle$ of ${\mathbb{F}}_{q^{n}}$ when $\textup{char}(F)=p\nmid n$. Since $\sum_{i=0}^{n-1}y^{q^{i}}=0$, the roots of $t^{q}-t-y$ lie in the (potentially larger) ${\mathbb{F}}_{q}$-subspace $\{x\in E\mid\textup{Tr}^{E}_{F}(x)=0\}$ of $E$ of codimension 1.
In the case when $n_{p}=p$, i.e. $p\mid n$ and $p^{2}\nmid n$, the element $z$ in Corollary 4.1 can be chosen to lie in the fixed field of $\sigma^{p}$, and to have a specific minimal polynomial.
Corollary 4.4.
Let $E/F$ be a cyclic field extension of degree $n$ where $n_{p}=\textup{char}(F)=p$. Suppose that $\textup{Gal}(E/F)=\langle\sigma\rangle$ where $\sigma\colon E\to E$ is the $p$th power map $\sigma(x)=x^{p}$. If $y\in E$ satisfies $\textup{Tr}^{E}_{F}(y)=\sum_{i=0}^{n-1}y^{q^{i}}=0$, and $\zeta$ satisfies $\zeta^{p}-\zeta^{p-1}+1=0$. Then $x^{p}-x-y=0$ where $x=(n/p)^{-1}\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}\zeta^{p^{j}}y^{p^{i}}$.
Proof.
It follows from $n_{p}=p$, that $n_{p^{\prime}}=n/p$ is coprime to $p$. Let $K$ be the fixed field of $\sigma^{p}$. Then $K$ contains an element $\zeta$ satisfying $\zeta^{p}-\zeta^{p-1}+1=0$ and $\textup{Tr}^{L}_{F}(\zeta)=1$ by Lemma 2.2. Therefore $\textup{Tr}^{E}_{F}(\zeta)=n/p$ and $\textup{Tr}^{E}_{F}(z)=1$ where $z=(n/p)^{-1}\zeta$. By Theorem 2.1(a), $x=(n/p)^{-1}\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}\zeta^{p^{j}}y^{p^{i}}$ satisfies $x^{p}-x-y=0$. ∎
Let $p,r$ be distinct primes and set $q=p^{f}$. Denote by $\textup{ord}_{r}(q)$ the smallest positive integer $e$ such that $q^{e}\equiv 1\pmod{r}$ and note that $e\mid(r-1)$. If $r\geqslant 2$, then the cyclotomic polynomial $\Phi_{r}(t)$ of degree $\phi(r)$ factors over ${\mathbb{F}}_{p}$ into irreducible polynomials each of degree $\textup{ord}_{r}(p)$ by [5, Theorem 2.47].
Corollary 4.1 shows that $z$ may be chosen to lie in a subfield of order $p^{n_{p}}$. Hence $z$ is a $(p^{n_{p}}-1)$th root of unity. Lemma 4.5 proves that, in many cases, $z$ may be chosen to have prime order $r\neq p$. This choice of $z$ may increase the period of the sequence $x_{0},x_{1},\dots$.
@@
Let $p,r$ be unequal primes where $\textup{ord}_{r}(p)\kern-1.0pt=\kern-1.0pte$. Let $E\kern-1.0pt=\kern-1.0pt{\mathbb{F}}_{p^{n}}$, $K\kern-1.0pt=\kern-1.0pt{\mathbb{F}}_{p^{e}}$, $F\kern-1.0pt=\kern-1.0pt{\mathbb{F}}_{p}$ where $e\mid n$. There exists a primitive $r$th root of unity $\zeta\in K$ for which $\tau\coloneq\textup{Tr}^{K}_{F}(\zeta)\neq 0$. If $p\nmid\frac{n}{e}$, and $y\in E$ satisfies $\textup{Tr}^{E}_{F}(y)=0$, then $x=(\frac{n\tau}{e})^{-1}\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}\zeta^{p^{j}}\right)y^{p^{i}}$ satisfies $x^{p}-x-y=0$. Finally, the sequence $(\sum_{j=0}^{i-1}\zeta^{p^{j}})_{i\geqslant 0}$ is periodic with period $ep$ where $n_{p}\mid e$.
Proof.
First, $r$ divides $p^{d}-1$ if and only if $e\mid d$. As $e\mid n$, all primitive $r$th roots of unity lie in $K$, and $F\leqslant K\leqslant E$ holds. The cyclotomic polynomial $\Phi_{r}(t)=t^{r-1}+t^{r-2}+\cdots+t+1$ factors over $F$ as a product of $(r-1)/e$ irreducible polynomials each of degree $e$, so the coefficient of $t^{e-1}$ must be nonzero in at least one irreducible. Hence some $r$th root of unity $\zeta\in K$ has $\tau\coloneq\textup{Tr}^{K}_{F}(\zeta)\neq 0$. Thus $\textup{Tr}^{E}_{K}(\zeta)=\frac{n}{e}\zeta$, so that $\textup{Tr}^{E}_{F}(\zeta)=\textup{Tr}^{K}_{F}(\textup{Tr}^{E}_{K}(\zeta))=\frac{n}{e}\textup{Tr}^{K}_{F}(\zeta)=\frac{n\tau}{e}\neq 0$. Therefore $z=(\frac{n\tau}{e})^{-1}\zeta\in K^{\times}$ satisfies $\textup{Tr}^{E}_{F}(z)=1$. The formula for $x=R(y,z)$ and the periodicity claim follow from Theorem 3.1(a,c). ∎
Corollary 4.6.
Let $p$ be a prime where $p\equiv 2\pmod{3}$. Let $E={\mathbb{F}}_{p^{n}}$ and $F={\mathbb{F}}_{p}$ where $n$ is even and let $\omega\in E$ satisfy $\omega^{2}+\omega+1=0$. If $p\nmid(n/2)$, and $y\in E$ satisfies $\textup{Tr}^{E}_{F}(y)=0$, then the sum $R(y)=(\frac{n}{2})^{-1}\sum_{i=0}^{n-1}\left(\lfloor\frac{i}{2}\rfloor-r\omega\right)y^{p^{i}}$ satisfies $R(y)^{p}-R(y)-y=0$ where $r\in\{0,1\}$ and $i\equiv r\pmod{2}$. Finally, the sequence $x_{0},x_{1},\dots$ is periodic of period $2p$ where $x_{i}=(\frac{n}{2})^{-1}\left(\lfloor\frac{i}{2}\rfloor-r\omega\right)$.
Proof.
The condition $p\equiv 2\pmod{3}$ is equivalent to $\textup{ord}_{3}(p)=2$. Thus we may use Lemma 4.5 with $r=3$ and $\zeta$ replaced with $\omega$. However, $\omega^{p^{k}}$ is $\omega$ if $k$ is even, and $\omega^{2}$ if $k$ is odd. If $i=0,1,2,3,4,5,6$, then $s_{i}=\sum_{k=0}^{i-1}\omega^{p^{k}}$ equals 00, $\omega$, $\omega+\omega^{2}=-1$, $-1+\omega$, $-2$, $-2+\omega$, $-3$, etc. Therefore $s_{i}=-(i-r)/2+r\omega$ where $r\equiv i\pmod{2}\in\{0,1\}$ and $r\in\{0,1\}$. The values of $x_{i}=(\frac{n}{2})^{-1}\left(\frac{i-r}{2}-r\omega\right)$ are periodic of period $2p$ by Theorem 3.1(c). ∎
Remark 4.7.
If $n_{p}$ is bounded, then Corollary 4.1 gives a closed-form expression $R(y,z)=\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}z^{q^{j}}\right)y^{q^{i}}$ for a root of a reducible polynomial $t^{q}-t-y$ over ${\mathbb{F}}_{q^{n}}$. Further, the element $z$ may be chosen to have degree $n_{p}$, the $p$-part of $n$, and the sequence $\left(\sum_{j=0}^{i-1}z^{q^{j}}\right)_{i\geqslant 0}$ has period $pn_{p}$. If $n_{p}$ is unbounded, then $R(y,z)$ is not a closed-form expression as it involves infinitely many choices for an element $z$ satisfying $\textup{Tr}^{E}_{F}(z)=1$.
The following result gives infinitely many ‘quadratic formulas’ over the finite field ${\mathbb{F}}_{2^{n}}$.
Lemma 4.8.
Let $y\in{\mathbb{F}}_{2^{n}}$ where $\sum_{i=0}^{n-1}y^{2^{i}}=0$. Then $x=\sum_{i=0}^{n-1}x_{i}y^{2^{i}}$ is a root of $t^{2}+t+y$ where $x_{i}=\sum_{j=0}^{i-1}z^{2^{j}}$ and $z\in{\mathbb{F}}_{2^{n}}$ has degree $n_{2}$. Write $z=(n/n_{2})^{-1}\zeta$. Table 4.8 lists choices for $\zeta$ and the minimal polynomial $m_{\zeta/{\mathbb{F}}_{2}}(t)$ of $\zeta$, as a function of the value of $n_{2}$.
@@
Set $E={\mathbb{F}}_{2^{n}}$ and $F={\mathbb{F}}_{2}$. The polynomials in the last line of Table 4.8 are irreducible over $F$ (and hence can equal $m_{\zeta/F}(t)$), and $\zeta$ has order $12^{n_{2}}-$ 1. By Theorem 3.1, $x=R(y,z)$ is a root of $t^{2}+t+y$ where $z\in E$ is an element of degree $e=n_{2}$ satisfying $\textup{Tr}^{E}_{F}(z)=1$. Let $K$ be the subfield of $E$ with $|K:F|=n_{2}$ and $|E:K|=n_{2^{\prime}}$. We choose an element $\zeta\in K$ with $K={\mathbb{F}}_{2}(\zeta)$ and $\textup{Tr}^{K}_{F}(\zeta)=1$, and set $z=(n_{2^{\prime}})^{-1}\zeta$. If $s_{i}=\sum_{j=0}^{i-1}\zeta^{2^{j}}$, then $s_{0}=s_{2e}=0$ and $s_{i+2e}=s_{i}$, so it suffices to list $x_{i}=(n_{2^{\prime}})^{-1}s_{i}$ for $0\leqslant i<2e$. Indeed, since $s_{e}=1$ and $s_{i+e}=1+s_{i}$, it suffices to list $x_{i}$ for $0\leqslant i<e$.
@@
(b) Suppose that $n_{2}=2$ or $n\equiv 2\pmod{4}$. Set $\zeta=\omega$ or set $p=2$ into Corollary 4.6. Since $2\nmid(n/2)$ the value of $x_{i}=(\frac{n}{2})^{-1}\left(\lfloor\frac{i}{2}\rfloor-r\omega\right)$ is $0,\omega,1,\omega^{2}$ when $i\equiv 0,1,2,3\pmod{4}$. The formulas for $x$ in cases (a) and (b) were given previously by Lahtonen [3].
(c) Suppose that $n_{2}=4$, i.e. $n\equiv 4\pmod{8}$. Then $\zeta=\alpha\in{\mathbb{F}}_{16}$ has order $12^{4}-$ 1 and satisfies $\textup{Tr}^{K}_{F}(\zeta)=1$. Hence $s_{i}=0,\alpha,\alpha^{13},\alpha^{6},1,\alpha^{12},\alpha^{7},\alpha^{8}$ for $i\equiv 0,1,\dots,7\pmod{8}$.
(d) Suppose that $n_{2}=8$, i.e. $n\equiv 8\pmod{16}$. Then $\zeta=\beta\in{\mathbb{F}}_{256}$ has order $12^{8}-$ 1 and satisfies $\textup{Tr}^{K}_{F}(\zeta)=1$. Hence $s_{i}=0,\beta,\beta^{100},\beta^{189},\beta^{29},\beta^{60},\beta^{154},\beta^{177}$ for $i\equiv 0,1,\dots,7\pmod{16}$. Also $s_{i+8}=1+s_{i}$ for $i\equiv 0,1,\dots,7\pmod{16}$.
(e) Suppose that $n_{2}=16$, i.e. $n\equiv 16\pmod{32}$. Then $\zeta=\gamma\in{\mathbb{F}}_{2^{16}}$ has order $12^{16}-$ 1 and $\textup{Tr}^{K}_{F}(\zeta)=1$. Hence $s_{i}=0,\gamma,\gamma^{64409},\gamma^{48754},\gamma^{27742},\gamma^{48469},\gamma^{1146},\gamma^{22404},\gamma^{64313},\gamma^{47682},$ $\gamma^{63219},\gamma^{45929},\gamma^{55680},\gamma^{46875},\gamma^{7495},\gamma^{32204}$ and $s_{i+16}=1+s_{i}$ for $i\equiv 0,1,\dots,15\pmod{32}$. ∎

### v2

@@ 3 Periodicity and Hilbert’s additive Theorem 90
We prove in Theorem 1.2 that when $\textup{char}(F)=p>0$, the element $x$ in Theorem 1.1 has an intrinsic periodicity. This observation is used in Corollary 4.1 to give a polynomial formula for the roots of (reducible) Artin-Schreier polynomials.
Proof of Theorem 1.2.
(a) First, we show that the fixed subfield $K$ of $\sigma^{n_{p}}$ contains an element $z$ satisfying $\textup{Tr}^{E}_{F}(z)=1$. Since $n_{p^{\prime}}=n/n_{p}\neq 0$ in $F$ and the map $\textup{Tr}^{K}_{F}\colon K\to F$ is surjective, there exists an element $z\in K$ such that $\textup{Tr}^{K}_{F}(z)=(n_{p^{\prime}})^{-1}$. Therefore $\textup{Tr}^{E}_{K}(z)=|E:K|z=n_{p^{\prime}}z$ and hence $\textup{Tr}^{E}_{F}(z)=\textup{Tr}^{K}_{F}(\textup{Tr}^{E}_{K}(z))=n_{p^{\prime}}\textup{Tr}^{K}_{F}(z)=1$. Next, suppose that $L=F(z)$ and $\textup{Tr}^{E}_{F}(z)=1$. Then $1=\textup{Tr}^{L}_{F}(\textup{Tr}^{E}_{L}(z))=|E:L|\textup{Tr}^{L}_{F}(z)$. Therefore $p\nmid|E:L|$, so $n_{p}=|K:F|$ divides $e=|L:F|$ as claimed. In summary, the minimal polynomial over $F$ of an element $z$ with $\textup{Tr}^{E}_{F}(z)=1$ has degree a multiple of $n_{p}$, and degree precisely $n_{p}$ if $z$ lies in $K\setminus K_{0}$ where $K_{0}$ is the unique (maximal) subfield of $K$ with $F\leqslant K_{0}<K$ and $|K:K_{0}|=p$.
(b) Let $x_{i}=\sum_{j=0}^{i-1}\sigma^{j}z$ for $0\leqslant i<n$ where $\textup{Tr}^{E}_{F}(z)=1$. Let $L=F(z)$ and $e=|L:F|$. Hence $x_{e}=\textup{Tr}^{L}_{F}(z)\neq 0$ by part (b) and $\sigma^{j}z=z$ implies $e\mid j$. Suppose that $x_{k}=0$ where $0\leqslant k<pe$ and write $k=ie+j$ where $0\leqslant i<p$ and $0\leqslant j<e$. The equation $x_{ie+j}=ix_{e}+x_{j}$ shows that $x_{0}=x_{pe}=0$. Thus $x_{k}=ix_{e}+x_{j}=0$ and $x_{e}\in F$ gives
$x_{j+1}-z=\sigma x_{j}=\sigma(-ix_{e})=-ix_{e}=x_{j}.$
4 Formulas for the roots of Artin-Schreier polynomials
In this section we assume that $F$ has prime characteristic $p$. Set $q=p^{f}$. Two examples with $\textup{Gal}(E/F)=\langle\sigma\rangle$ of order $n$ and $\sigma(x)=x^{q}$ for $x\in E$ are the finite fields $E={\mathbb{F}}_{q^{n}}$, $F={\mathbb{F}}_{q}$, and the function fields $E={\mathbb{F}}_{q^{n}}(u)$, $F={\mathbb{F}}_{q}(u)$. We shall give a number of different applications of Theorem 1.2 depending on the value of $n_{p}$. Observe first that Theorem 1.1 gives a polynomial formula for $x$ when $\sigma(x)=x^{q}$ as noted below.
Corollary 4.1.
Suppose that $\textup{char}(F)=p$, $q=p^{f}$, and $E/F$ is a cyclic extension of degree $n$ where $\textup{Gal}(E/F)=\langle\sigma\rangle$ and $\sigma(x)=x^{q}$. If $y\in E$ and $\textup{Tr}^{E}_{F}(y)=0$, then $x=\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}z^{q^{j}}y^{q^{i}}$ satisfies $x^{q}-x-y=0$.
When $n$ is coprime to $p$, i.e. $n_{p}=1$, the formula for $x$ in Corollary 4.1 simplifies.
Corollary 4.2.
Suppose that ${\mathbb{F}}_{q}\leqslant F$, $q=p^{f}$, and $E/F$ is a cyclic extension of degree $n$ where $\textup{Gal}(E/F)=\langle\sigma\rangle$ and $\sigma(x)=x^{q}$. If $y\in E$, $\textup{Tr}^{E}_{F}(y)=0$ and $n_{p}=1$, then $t^{q}-t-y=\prod_{u\in{\mathbb{F}}_{q}}(t-x-u)$ where $x=\sum_{i=0}^{n-1}\frac{i}{n}y^{q^{i}}$.
Proof.
By Theorem 1.2(b), we may choose $z$ to lie in $F$ and so $1=\textup{Tr}^{E}_{F}(z)=nz$. Hence $z=\frac{1}{n}$ and $x=R(y,z)=\sum_{i=0}^{n-1}\frac{i}{n}y^{q^{i}}$ satisfies $x^{q}-x-y=0$ by Corollary 4.1. Therefore $x+u$ is a root of $t^{q}-t-y$ for each $u\in{\mathbb{F}}_{q}\leqslant F$, and the claimed factorization follows.∎
@@ 4 Formulas for the roots of Artin-Schreier polynomials
When $\textup{char}(F)=p\nmid n$, Corollary 4.2 implies that all the roots of $t^{q}-t-y$ lie in the ${\mathbb{F}}_{p}$-subspace $\langle y,y^{q},\dots,y^{q^{n-1}}\rangle$ of ${\mathbb{F}}_{q^{n}}$. Since $\sum_{i=0}^{n-1}y^{q^{i}}=0$, the roots of $t^{q}-t-y$ lie in the (potentially larger) ${\mathbb{F}}_{q}$-subspace $\{w\in E\mid\textup{Tr}^{E}_{F}(w)=0\}$ of $E$ of codimension 1.
When $n_{p}=p$, i.e. $p\mid n$ and $p^{2}\nmid n$, the element $z$ in Corollary 4.1 can be chosen to lie in the fixed field of $\sigma^{p}$, and to have a specific minimal polynomial.
@@
Let $E/F$ be a cyclic field extension of degree $n$ where $n_{p}=\textup{char}(F)=p$. Suppose that $\textup{Gal}(E/F)=\langle\sigma\rangle$ where $\sigma\colon E\to E$ is the $p$th power map $\sigma(x)=x^{p}$. If $y\in E$ satisfies $\textup{Tr}^{E}_{F}(y)=\sum_{i=0}^{n-1}y^{p^{i}}=0$, and $\zeta$ satisfies $\zeta^{p}-\zeta^{p-1}+1=0$. Then $x^{p}-x-y=0$ where $x=(n/p)^{-1}\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}\zeta^{p^{j}}y^{p^{i}}$ and $\zeta^{p^{p}-1}=1$.
Proof.
It follows from $n_{p}=p$, that $n_{p^{\prime}}=n/p$ is coprime to $p$. Let $K$ be the fixed field of $\sigma^{p}$. Then $K$ contains an element $\zeta$ satisfying $\zeta^{p}-\zeta^{p-1}+1=0$ and $\textup{Tr}^{L}_{F}(\zeta)=1$ by Lemma 2.2. Therefore $\textup{Tr}^{E}_{F}(\zeta)=n/p$ and $\textup{Tr}^{E}_{F}(z)=1$ where $z=(n/p)^{-1}\zeta$. By Theorem 1.1, $x=(n/p)^{-1}\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}\zeta^{p^{j}}y^{p^{i}}$ satisfies $x^{p}-x-y=0$. Finally, $\sigma(\zeta)=\zeta^{p}$ implies that $\zeta=\sigma^{p}(\zeta)=\zeta^{p^{p}}$, so that $\zeta^{p^{p}-1}=1$ as $\zeta\neq 1$. ∎
Let $p,r$ be distinct primes. Denote by $\textup{ord}_{r}(p)$ the smallest positive integer $e$ such that $p^{e}\equiv 1\pmod{r}$ and note that $e\mid(r-1)$. The cyclotomic polynomial $\Phi_{r}(t)=\sum_{i=0}^{r-1}t^{i}$ of degree $r-1$ factors over ${\mathbb{F}}_{p}$ as a product of $(r-1)/e$ irreducible polynomials each of degree $e=\textup{ord}_{r}(p)$ by [6, Theorem 2.47].
Theorem 1.2 shows that $z$ may be chosen to lie in a subfield of order $p^{f}$ where $f={n_{p}}$. Hence $z^{p^{f}-1}=1$. Lemma 4.5 proves that, in many cases, $z$ may be chosen to have prime order $r\neq p$. This choice of $z$, however, may increase the period of the sequence $(x_{i})_{i\geqslant 0}$.
Lemma 4.5.
Let $p,r$ be unequal primes. Let $e\coloneq\textup{ord}_{r}(p)$ divide $n$. Set $E={\mathbb{F}}_{p^{n}}$, $K={\mathbb{F}}_{p^{e}}$ and $F={\mathbb{F}}_{p}$. There exists a primitive $r$th root of unity $\zeta\in K$ for which $\tau\coloneq\textup{Tr}^{K}_{F}(\zeta)\neq 0$. If $p\nmid\frac{n}{e}$ and $y\in E$ has $\textup{Tr}^{E}_{F}(y)=0$, then $x=(\frac{n\tau}{e})^{-1}\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}\zeta^{p^{j}}\right)y^{p^{i}}$ satisfies $x^{p}-x-y=0$. Finally, the sequence $(\sum_{j=0}^{i-1}\zeta^{p^{j}})_{i\geqslant 0}$ is periodic with period $ep$ where $n_{p}\mid e$.
Proof.
First, $r$ divides $p^{d}-1$ if and only if $e\mid d$. As $e\mid n$, all primitive $r$th roots of unity lie in $K$, and $F\leqslant K\leqslant E$ holds. The cyclotomic polynomial $\Phi_{r}(t)$ factors over $F$ as a product of $(r-1)/e$ irreducible polynomials each of degree $e$, so the coefficient of $t^{e-1}$ must be nonzero in at least one irreducible. Hence some $r$th root of unity $\zeta\in K$ has $\tau\coloneq\textup{Tr}^{K}_{F}(\zeta)\neq 0$. Thus $\textup{Tr}^{E}_{K}(\zeta)=\frac{n}{e}\zeta$ and $\textup{Tr}^{E}_{F}(\zeta)=\textup{Tr}^{K}_{F}(\textup{Tr}^{E}_{K}(\zeta))=\frac{n}{e}\textup{Tr}^{K}_{F}(\zeta)=\frac{n\tau}{e}\neq 0$. Therefore $z=(\frac{n\tau}{e})^{-1}\zeta\in K^{\times}$ satisfies $\textup{Tr}^{E}_{F}(z)=1$. The formula for $x=R(y,z)$ follows from Theorem 1.1 and the periodicity $ep$ follows from Theorem 1.2(b). ∎
@@
Let $p$ be a prime where $p\equiv 2\pmod{3}$. Let $E={\mathbb{F}}_{p^{n}}$ and $F={\mathbb{F}}_{p}$ where $n$ is even and let $\omega\in E$ satisfy $\omega^{2}+\omega+1=0$. If $p\nmid(n/2)$, and $y\in E$ satisfies $\textup{Tr}^{E}_{F}(y)=0$, then the sum $x=(\frac{n}{2})^{-1}\sum_{i=0}^{n-1}\left(\lfloor\frac{i}{2}\rfloor-r_{i}\omega\right)y^{p^{i}}$ satisfies $x^{p}-x-y=0$ where $r_{i}\in\{0,1\}$ and $i\equiv r_{i}\pmod{2}$. Finally, the sequence $x_{0},x_{1},\dots$ is periodic of period $2p$ where $x_{i}=(\frac{n}{2})^{-1}\left(\lfloor\frac{i}{2}\rfloor-r_{i}\omega\right)$ and $r_{i}=i\mod 2\in\{0,1\}$.
@@
Since $n$ is even and $p^{n}\equiv 1\pmod{3}$, $E$ contains a primitive cube root $\omega$ of 1. The condition $p\equiv 2\pmod{3}$ is equivalent to $\textup{ord}_{3}(p)=2$. Thus we may use Lemma 4.5 with $r=3$ and $\zeta=\omega$. However, $\omega^{p^{k}}$ is $\omega$ if $k$ is even, and $\omega^{2}$ if $k$ is odd. If $i=0,1,2,3,4,5,6$, then $s_{i}=\sum_{k=0}^{i-1}\omega^{p^{k}}$ equals 00, $\omega$, $\omega+\omega^{2}=-1$, $-1+\omega$, $-2$, $-2+\omega$, $-3$, etc. Therefore $s_{i}=-(i-r_{i})/2+r_{i}\omega$ where $r_{i}\equiv i\pmod{2}\in\{0,1\}$ and $r_{i}\in\{0,1\}$. The values of $x_{i}=(\frac{n}{2})^{-1}\left(\frac{i-r_{i}}{2}-r_{i}\omega\right)$ are periodic of period $2p$ by Theorem 1.2(b). Finally, $\frac{i-r_{i}}{2}=\lfloor\frac{i}{2}\rfloor$. ∎
The following result gives different quadratic ‘formulas’ over the finite field ${\mathbb{F}}_{2^{n}}$ that depend on the 2-part $n_{2}$ of $n$ and the choice of an element $z\in{\mathbb{F}}_{2^{n}}$ with absolute trace 1. The condition $n_{2}=2^{k}$ is equivalent to $n\equiv 2^{k}\pmod{2^{k+1}}$.
@@
Let $y\in{\mathbb{F}}_{2^{n}}$ where $\sum_{i=0}^{n-1}y^{2^{i}}=0$. Then $x=\sum_{i=0}^{n-1}x_{i}y^{2^{i}}$ is a root of $t^{2}+t+y$ where $x_{i}=\sum_{j=0}^{i-1}z^{2^{j}}$ and $z\in{\mathbb{F}}_{2^{n}}$ has absolute trace $\sum_{i=0}^{n-1}z^{2^{i}}=1$ and has degree $n_{2}$. Write $z=(n/n_{2})^{-1}\zeta$. Table 4.7 lists choices for $\zeta$ and the minimal polynomial $m_{\zeta/{\mathbb{F}}_{2}}(t)$ of $\zeta$, as a function of the value of $n_{2}$.
@@
Set $E={\mathbb{F}}_{2^{n}}$ and $F={\mathbb{F}}_{2}$. The polynomials in the last line of Table 4.7 are irreducible over $F$ (and hence can equal $m_{\zeta/F}(t)$), and $\zeta$ has order $12^{n_{2}}-$ 1. By Theorem 1.1, $x=R(y,z)$ is a root of $t^{2}+t+y$ where $z\in E$ is an element of degree $e=n_{2}$ satisfying $\textup{Tr}^{E}_{F}(z)=1$. Let $K$ be the subfield of $E$ with $|K:F|=n_{2}$ and $|E:K|=n_{2^{\prime}}$. We choose an element $\zeta\in K$ with $K={\mathbb{F}}_{2}(\zeta)$ and $\textup{Tr}^{K}_{F}(\zeta)=1$, and set $z=(n_{2^{\prime}})^{-1}\zeta$. Let $s_{i}=\sum_{j=0}^{i-1}\zeta^{2^{j}}$. Then $s_{0}=s_{2e}=0$ and $s_{i+2e}=s_{i}$, so it suffices to list $x_{i}=(n_{2^{\prime}})^{-1}s_{i}$ for $0\leqslant i<2e$. Indeed, since $s_{e}=1$ and $s_{i+e}=1+s_{i}$, it suffices to list $x_{i}$ for $0\leqslant i<e$.
@@
(b) Suppose that $n_{2}=2$. Set $\zeta=\omega$ or set $p=2$ into Corollary 4.6. Since $2\nmid(n/2)$ the value of $x_{i}=(\frac{n}{2})^{-1}\left(\lfloor\frac{i}{2}\rfloor-r_{i}\omega\right)$ is $0,\omega,1,\omega^{2}$ when $i\equiv 0,1,2,3\pmod{4}$. (Recall that $r_{i}=i-2\lfloor\frac{i}{2}\rfloor\in\{0,1\}$.) Equivalent formulas for $x$ in cases (a) and (b) were given previously by Chen [1, Theorems 1, 2] and by Lahtonen [4].
(c) Suppose that $n_{2}=4$. Then $\zeta=\alpha\in{\mathbb{F}}_{16}$ has order $12^{4}-$ 1 and $\textup{Tr}^{K}_{F}(\zeta)=1$. Further, the values of $s_{i}=\sum_{j=0}^{i-1}\zeta^{2^{i}}$ are $0,\alpha,\alpha^{13},\alpha^{6},1,\alpha^{12},\alpha^{7},\alpha^{8}$ for $i\equiv 0,1,\dots,7\pmod{8}$.
(d) Suppose that $n_{2}=8$. Then $\zeta=\beta\in{\mathbb{F}}_{256}$ has order $12^{8}-$ 1 and satisfies $\textup{Tr}^{K}_{F}(\zeta)=1$. Furthermore, the values of $s_{i}=\sum_{j=0}^{i-1}\zeta^{2^{i}}$ are $0,\beta,\beta^{100},\beta^{189},\beta^{29},\beta^{60},\beta^{154},\beta^{177}$ for $i\equiv 0,1,\dots,7\pmod{16}$. Also $s_{i+8}=1+s_{i}$ for $i\equiv 0,1,\dots,7\pmod{16}$.
(e) Suppose that $n_{2}=16$. Then $\zeta=\gamma\in{\mathbb{F}}_{2^{16}}$ has order $12^{16}-$ 1 and $\textup{Tr}^{K}_{F}(\zeta)=1$. In addition, $s_{i}=\sum_{j=0}^{i-1}\zeta^{2^{j}}$ equals $0,\gamma,\gamma^{64409},\gamma^{48754},\gamma^{27742},\gamma^{48469},\gamma^{1146},\gamma^{22404},\gamma^{64313},\gamma^{47682},$ $\gamma^{63219},\gamma^{45929},\gamma^{55680},\gamma^{46875},\gamma^{7495},\gamma^{32204}$ and $s_{i+16}=1+s_{i}$ for $i\equiv 0,1,\dots,15\pmod{32}$. ∎